# Quasinormal Modes and Greybody Factors of Charged Symmergent Black Hole

### v1

@@ IV Perturbations and Quasinormal modes
In this section, we focus on the analysis of massless scalar and vector perturbations within the framework of the black hole metric defined in equation (12). For the sake of simplicity, we assume that the test field, whether a scalar field or a vector field, exerts a negligible influence on the black hole spacetime, meaning that the back reaction of the field on the spacetime geometry is insignificant lopez2020 ; Chandrasekhar:1985kt . This assumption allows us to treat the perturbations as propagating on a fixed background metric without needing to consider the complex, coupled dynamics that would arise if the back reaction were significant.
@@
Our analysis specifically focuses on axial perturbations. To describe these perturbations, we express the perturbed metric as shown in the following way lopez2020
@@
IV.1 Massless Scalar Perturbation
@@
where $l$ and $m$ represent the polar and azimuthal indices, respectively sommerfeld ; Chandrasekhar:1985kt . The function $\psi_{s}(t,r)$ is the radial time-dependent wave function and the index “s” stands for scalar.
@@
Here, $l$ corresponds to the multipole moment of the black hole’s QNMs, and $\omega$ represents the frequency of these modes, which is determined from $\partial_{t}^{2}\psi_{s}(t,r_{*})=-\omega^{2}\psi_{s}(t,r_{*})$. The wavefunction $\psi(r_{*})_{s}$ thus describes behaviour of the scalar QNMs of the CSBH Chandrasekhar:1985kt ; lopez2020 .
@@
IV.2 Massless Vector or Electromagnetic Perturbation
@@
While these vector perturbations can be identified with the electromagnetic field, the broader symmergent gravity framework predicts the existence of additional massless particles, which may not interact with known particles through non-gravitational forces. These could include massless scalar fields or massless (dark) photons, which could significantly influence gravitational wave emissions. The presence of such particles in the symmergent particle spectrum could lead to observable deviations in the dynamics of gravitational waves and other astrophysical phenomena dad2019 ; dad2023 ; dad2021 .
IV.3 Behaviour of the Perturbation Potentials
@@
The analysis of the perturbation potential provides valuable insights into the nature of the QNMs for both scalar and electromagnetic perturbations. The variations in the potential with respect to parameters such as the multipole moment $l$, the model parameter $\alpha$, the coupling parameter $c_{\rm O}$, and the black hole charge $Q$ reveal how sensitive the QNMs are to these factors. This understanding is crucial for predicting the stability and dynamical response of black holes to perturbations, particularly in the context of different gravity models or in the presence of additional fields.
IV.4 Time-domain profile and quasinormal modes
To begin, let us first examine the time-domain profile of the scalar field before delving into the determination of QNMs. In the preceding section, we undertook extensive numerical computations to characterize the QNMs, thoroughly analyzing how these modes depend on the model parameters $Q$, $\alpha$, and $c_{\rm O}$. Our approach involved solving the perturbation equations in the frequency domain, which provided insights into the oscillatory behavior and damping times of the QNMs. However, frequency-domain analysis alone does not capture the full temporal evolution of the perturbations. Thus, in the subsequent section, our focus shifts towards understanding the time-domain behavior of both scalar and electromagnetic perturbations. This involves generating temporal profiles that illustrate how these fields evolve with time after an initial disturbance. To achieve this, we employ a time-domain integration framework based on the methodology proposed by Gundlach et al. gundlach , which is well-suited for capturing the full dynamical evolution of perturbations. By applying this technique, we can observe the decay patterns and the late-time behavior of the perturbations, offering a complementary perspective to the frequency-domain analysis. This dual approach not only helps in verifying the consistency of our results but also provides a more comprehensive understanding of the dynamical properties of the perturbations in the given gravitational background.
@@ IV.4 Time-domain profile and quasinormal modes
IV.5 WKB method with Padé Approximation for Quasinormal modes
In this study, we aim to estimate the QNMs of the black holes by employing the Wentzel-Kramers-Brillouin (WKB) method, a widely recognized approximation technique in the field of black hole perturbation theory. Initially introduced by Schutz and Will Schutz , the WKB method provides a first-order approximation for calculating QNMs. However, despite its utility, the method is known to exhibit certain limitations, including a higher degree of error in some cases. To address these limitations, researchers have developed higher-order WKB approximations, significantly improving the accuracy of QNM calculations Will_wkb ; Konoplya_wkb ; Maty_wkb . Our work utilizes the advancements suggested by Ref. Maty_wkb , where Padé approximations are incorporated into the WKB method, yielding more precise results.
Building on these improvements, we apply the Padé-averaged 6th-order WKB approximation technique to estimate the QNMs of black holes. The incorporation of Padé averaging has been shown to enhance the precision of QNM estimates, as demonstrated in prior studies such as Ref. Konoplya_wkb . By using this refined approach, we seek to achieve more accurate and reliable estimates of QNMs, which will be compared with earlier results to validate our findings. This comparative analysis is critical in evaluating the robustness of the WKB method and advancing our understanding of black hole dynamics.
In this subsection, we present QNM values computed using the Padé-averaged 6th-order WKB approximation, along with error estimates in Tables 1 and 2. The 3rd and 4th columns of these tables show the errors associated with the WKB approximation, specifically the root mean square (rms) error, denoted as $\vartriangle_{rms}$, and the error term $\Delta_{6}$, defined as Konoplya_wkb
@@ IV.5 WKB method with Padé Approximation for Quasinormal modes
These results highlight that the CSBH model exhibits behavior distinct from other well-known charged black hole solutions, such as the Reissner-Nordström black hole. In addition, the impact of the symmergent parameters on the QNM spectrum is distinct from that of the black hole’s charge, further emphasizing the uniqueness of this model. The study of QNMs in charged black holes, and their dependence on parameters such as $Q$, continues to be an important area of research for both theoretical and observational astrophysics Yang:2022ifo ; Daghigh:2008jz ; Zhidenko:2003wq ; Zhidenko:2005mv .
These findings also differ from results obtained for wormhole configurations, as studied in Ovgun188 ; gogoi_wormhole , demonstrating that the symmergent model’s predictions for QNMs are distinct even when compared to other exotic spacetime geometries. This underscores the importance of QNMs as a tool for probing the underlying nature of black holes and the influence of additional parameters such as charge and symmergent contributions on their dynamical properties.
@@
V Greybody factors
The concept of greybody factors originates from Hawking’s groundbreaking discovery in 1975, which demonstrated that black holes are not completely black but emit radiation, now known as Hawking radiation Hawking:1975vcx . This emission occurs near the black hole’s horizon, but the radiation that reaches a distant observer is altered by a redshift factor, resulting in a modification of its spectrum. This distortion is known as the greybody factor, and it reflects the difference between the initial Hawking radiation and what an observer at infinity detects Singleton:2011vh ; Akhmedova:2008dz . Various methods have been developed to calculate greybody factors, including significant contributions by Maldacena et al. Maldacena:1996ix , Fernando Fernando:2004ay , and others Okyay2022 ; Ovgun188 ; Pantig:2022gih ; Yang:2022xxh ; Yang:2022ifo ; Panotopoulos:2018pvu ; Panotopoulos:2016wuu ; Rincon:2018ktz ; Ahmed:2016lou ; Javed:2022kzf ; Al-Badawi:2022aby .
V.1 Using the WKB approach
@@
which involves the maximum value of the effective potential $V_{0}$, its second derivative $V_{0}^{\prime\prime}$, and higher-order corrections $\Lambda_{i}$, extending up to the 6th order Schutz ; Will_wkb ; Konoplya_wkb ; Maty_wkb . The WKB method, while effective, is less accurate at low frequencies where reflection tends to be total, causing the greybody factors to approach zero. Nevertheless, this does not substantially affect the calculation of energy emission rates.
Given the robustness and broad applicability of the WKB method in various contexts, including the study of black hole perturbations and greybody factors, a detailed review is beyond the scope of this paper. For more in-depth discussions, we refer readers to comprehensive reviews in the literature Konoplya:2019hlu ; Konoplya:2011qq .
@@
V.2 Rigorous Bounds on Greybody Factors
@@
The method we employ is based on the elegant analytical approach originally introduced by Visser (1998) Visser:1998ke , which was subsequently developed further by Boonserm and Visser (2008) Boonserm:2008zg . This technique has been widely applied in various contexts, as explored by numerous researchers, including Boonserm et al. (2017, 2019) Boonserm:2017qcq , Yang et al. (2022) Yang:2022ifo , Gray et al. (2015) Gray:2015xig , Ngampitipan et al. (2012) Ngampitipan:2012dq , and others Chowdhury:2020bdi ; Miao:2017jtr ; Liu:2021xfs ; Barman:2019vst . These studies have demonstrated the robustness of this approach in determining greybody factor bounds across different gravitational systems.
@@
Using this effective potential, we proceed to derive the rigorous lower bound for the greybody factors, following the methodology laid out by Visser (1998) Visser:1998ke and Boonserm and Visser (2008) Boonserm:2008zg . The bound for the transmission coefficient, denoted as $T_{b}$, is given by:
@@
To account for the presence of the cosmological constant and the symmergent gravity parameters, we modify the boundary conditions in accordance with the work by Boonserm et al. (2019) Boonserm:2019mon . The modified bound is expressed as:
@@
where $C_{1}=\left(r_{H}R_{H}+r_{H}^{2}+R_{H}^{2}\right)$ and the parameters $c_{\rm O}$, $\alpha$, and $Q$ represent the quadratic curvature term, the vacuum energy parameter, and the charge of the black hole, respectively.
@@
VI Concluding Remarks
In this paper, we have studied charged black hole solutions in the Symmergent gravity Pulice:2023dqw by investigating the QNMs and greybody factors of the CSBHs , focusing on scalar and electromagnetic perturbations. Our analyses reveal that the symmergent parameters $\alpha$, and the quadratic curvature term $c_{\rm O}$ have significant impacts on the QNMs and greybody factors as well as the black hole charge $Q$ and the multipole moment $l$.
@@ Acknowledgements.
A. Ö., D. J. G. and B. P. would like to acknowledge networking support by the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers). The work of B. P. is supported by the Astrophysics Research Center of the Open University of Israel (ARCO) through The Israeli Ministry of Regional Cooperation. The work of B. P. is supported by Sabancı University, Faculty of Engineering and Natural Sciences.

### v2

@@ III Perturbations and Quasinormal modes
In this section, we focus on the analysis of massless scalar and vector perturbations within the framework of the black hole metric defined in equation (12). For the sake of simplicity, we assume that the test field, whether a scalar field or a vector field, exerts a negligible influence on the black hole spacetime, meaning that the back reaction of the field on the spacetime geometry is insignificant [117, 49]. This assumption allows us to treat the perturbations as propagating on a fixed background metric without needing to consider the complex, coupled dynamics that would arise if the back reaction were significant.
@@
Our analysis specifically focuses on axial perturbations. To describe these perturbations, we express the perturbed metric as shown in the following way [117]
@@
III.1 Massless Scalar Perturbation
@@
where $l$ and $m$ represent the polar and azimuthal indices, respectively [118, 49]. The function $\psi_{s}(t,r)$ is the radial time-dependent wave function and the index “s” stands for scalar.
@@
Here, $l$ corresponds to the multipole moment of the black hole’s QNMs, and $\omega$ represents the frequency of these modes, which is determined from $\partial_{t}^{2}\psi_{s}(t,r_{*})=-\omega^{2}\psi_{s}(t,r_{*})$. The wavefunction $\psi(r_{*})_{s}$ thus describes behaviour of the scalar QNMs of the CSBH [49, 117].
@@
III.2 Massless Vector or Electromagnetic Perturbation
@@
While these vector perturbations can be identified with the electromagnetic field, the broader symmergent gravity framework predicts the existence of additional massless particles, which may not interact with known particles through non-gravitational forces. These could include massless scalar fields or massless (dark) photons, which could significantly influence gravitational wave emissions. The presence of such particles in the symmergent particle spectrum could lead to observable deviations in the dynamics of gravitational waves and other astrophysical phenomena [23, 16, 22].
III.3 Behaviour of the Perturbation Potentials
@@
The analysis of the perturbation potential provides valuable insights into the nature of the QNMs for both scalar and electromagnetic perturbations. The variations in the potential with respect to parameters such as the multipole moment $l$, the model parameter $\alpha$, the coupling parameter $c_{\rm O}$, and the black hole charge $Q$ reveal how sensitive the QNMs are to these factors. This understanding is crucial for predicting the stability and dynamical response of black holes to perturbations, particularly in the context of different gravity models or in the presence of additional fields. One may note that with $Q=0$, one can easily obtain the corresponding potentials discussed in Ref. [35]. However, we have noticed that the charge of the black hole i.e. $Q$ has a noticeable impact on the QNMs spectrum as predicted by the variation of the corresponding potential. In the presence of $Q$, the peak value of the potential increases and the variation is not similar to that of the Reissner-Nordström black hole. Hence, the QNMs spectrum seems to be different from that of a standard Reissner–Nordström black hole.
III.4 Time-domain profile and quasinormal modes
To begin, let us first examine the time-domain profile of the scalar field before delving into the determination of QNMs. In the preceding section, we undertook extensive numerical computations to characterize the QNMs, thoroughly analyzing how these modes depend on the model parameters $Q$, $\alpha$, and $c_{\rm O}$. Our approach involved solving the perturbation equations in the frequency domain, which provided insights into the oscillatory behavior and damping times of the QNMs. However, frequency-domain analysis alone does not capture the full temporal evolution of the perturbations. Thus, in the subsequent section, our focus shifts towards understanding the time-domain behavior of both scalar and electromagnetic perturbations. This involves generating temporal profiles that illustrate how these fields evolve with time after an initial disturbance. To achieve this, we employ a time-domain integration framework based on the methodology proposed by Gundlach et al. [127], which is well-suited for capturing the full dynamical evolution of perturbations. By applying this technique, we can observe the decay patterns and the late-time behavior of the perturbations, offering a complementary perspective to the frequency-domain analysis. This dual approach not only helps in verifying the consistency of our results but also provides a more comprehensive understanding of the dynamical properties of the perturbations in the given gravitational background.
@@ III.4 Time-domain profile and quasinormal modes
III.5 WKB method with Padé Approximation for Quasinormal modes
In this study, we aim to estimate the QNMs of the black holes by employing the Wentzel-Kramers-Brillouin (WKB) method, a widely recognized approximation technique in the field of black hole perturbation theory and investigate the validity of Hod’s conjecture. Initially introduced by Schutz and Will [123], the WKB method provides a first-order approximation for calculating QNMs. However, despite its utility, the method is known to exhibit certain limitations, including a higher degree of error in some cases. To address these limitations, researchers have developed higher-order WKB approximations, significantly improving the accuracy of QNM calculations [124, 125, 126]. Our work utilizes the advancements suggested by Ref. [126], where Padé approximations are incorporated into the WKB method, yielding more precise results.
Building on these improvements, we apply the Padé-averaged 6th-order WKB approximation technique to estimate the QNMs of black holes. The incorporation of Padé averaging has been shown to enhance the precision of QNM estimates, as demonstrated in prior studies such as Ref. [125]. By using this refined approach, we seek to achieve more accurate and reliable estimates of QNMs, which will be compared with earlier results to validate our findings. This comparative analysis is critical in evaluating the robustness of the WKB method and advancing our understanding of black hole dynamics.
In this subsection, we present QNM values computed using the Padé-averaged 6th-order WKB approximation, along with error estimates in Tables 1 and 2. The 3rd and 4th columns of these tables show the errors associated with the WKB approximation, specifically the root mean square (rms) error, denoted as $\vartriangle_{rms}$, and the error term $\Delta_{6}$, defined as [125]
@@ III.5 WKB method with Padé Approximation for Quasinormal modes
These results highlight that the CSBH model exhibits behavior distinct from other well-known charged black hole solutions, such as the Reissner-Nordström black hole. In addition, the impact of the symmergent parameters on the QNM spectrum is distinct from that of the black hole’s charge, further emphasizing the uniqueness of this model. The study of QNMs in charged black holes, and their dependence on parameters such as $Q$, continues to be an important area of research for both theoretical and observational astrophysics [95, 53, 98, 99].
These findings also differ from results obtained for wormhole configurations, as studied in [104, 105], demonstrating that the symmergent model’s predictions for QNMs are distinct even when compared to other exotic spacetime geometries. This underscores the importance of QNMs as a tool for probing the underlying nature of black holes and the influence of additional parameters such as charge and symmergent contributions on their dynamical properties.
@@
IV Greybody factors
The concept of greybody factors originates from Hawking’s groundbreaking discovery in 1975, which demonstrated that black holes are not completely black but emit radiation, now known as Hawking radiation [106]. This emission occurs near the black hole’s horizon, but the radiation that reaches a distant observer is altered by a redshift factor, resulting in a modification of its spectrum. This distortion is known as the greybody factor, and it reflects the difference between the initial Hawking radiation and what an observer at infinity detects [107, 108]. Various methods have been developed to calculate greybody factors, including significant contributions by Maldacena et al. [109], Fernando [110], and others [96, 104, 93, 111, 95, 112, 113, 101, 114, 115, 116].
IV.1 Using the WKB approach
@@
which involves the maximum value of the effective potential $V_{0}$, its second derivative $V_{0}^{\prime\prime}$, and higher-order corrections $\Lambda_{i}$, extending up to the 6th order [123, 124, 125, 126]. The WKB method, while effective, is less accurate at low frequencies where reflection tends to be total, causing the greybody factors to approach zero. Nevertheless, this does not substantially affect the calculation of energy emission rates.
Given the robustness and broad applicability of the WKB method in various contexts, including the study of black hole perturbations and greybody factors, a detailed review is beyond the scope of this paper. For more in-depth discussions, we refer readers to comprehensive reviews in the literature [128, 51].
@@
IV.2 Rigorous Bounds on Greybody Factors
@@
The method we employ is based on the elegant analytical approach originally introduced by Visser (1998) [129], which was subsequently developed further by Boonserm and Visser (2008) [130]. This technique has been widely applied in various contexts, as explored by numerous researchers, including Boonserm et al. (2017, 2019) [131], Yang et al. (2022) [95], Gray et al. (2015) [132], Ngampitipan et al. (2012) [133], and others [134, 135, 136, 137]. These studies have demonstrated the robustness of this approach in determining greybody factor bounds across different gravitational systems.
@@
Using this effective potential, we proceed to derive the rigorous lower bound for the greybody factors, following the methodology laid out by Visser (1998) [129] and Boonserm and Visser (2008) [130]. The bound for the transmission coefficient, denoted as $T_{b}$, is given by:
@@
To account for the presence of the cosmological constant and the symmergent gravity parameters, we modify the boundary conditions in accordance with the work by Boonserm et al. (2019) [138]. The modified bound is expressed as:
@@
where $C_{1}=\left(r_{H}R_{H}+r_{H}^{2}+R_{H}^{2}\right)$ and the parameters $c_{\rm O}$, $\alpha$, and $Q$ represent the quadratic curvature coupling parameter, the vacuum energy parameter, and the charge of the black hole, respectively.
@@
V Concluding Remarks
In this paper, we have studied charged black hole solutions in the Symmergent gravity [27] by investigating the QNMs and greybody factors of the CSBHs, focusing on scalar and electromagnetic perturbations. This work is an extension of our previous work [35]. Here, we have analysed the QNMs and greybody factors in the presence of charge $Q$. Our analyses reveal that the symmergent parameters $\alpha$, and the quadratic curvature coupling parameter $c_{\rm O}$ have significant impacts on the QNMs and greybody factors as well as the black hole charge $Q$ and the multipole moment $l$.
@@ Acknowledgements.
A. Ö., D. J. G. and B. P. would like to acknowledge networking support by the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers). The work of B. P. is supported by Sabancı University, Faculty of Engineering and Natural Sciences. The work of B. P. is supported by the Astrophysics Research Center of the Open University of Israel (ARCO) through The Israeli Ministry of Regional Cooperation.